# Self-similar decomposition of the hierarchical merger tree of dark matter halos

### v2

@@ Abstract
We show that the merger tree of dark matter halos is approximately self-similar by investigating the universality of the subhalo peak mass function (PMF) describing the mass distribution of progenitor halos. Using a set of cosmological simulations and identifying subhalos of different merger levels with hbt+, we verify that the level-1 subhalo PMF is close to universal across halo mass, redshift, and cosmology. This approximate self-similarity allows us to analytically derive the subhalo PMF for subhalos accreted at any level (i.e., for sub-sub…halos) through self-convolutions of the level-1 PMF, and the resulting model shows good agreement with simulation measurements. We further derive a number of analytical properties on the hierarchical origin of subhalos, including the level distribution, accretion rate at each level, initial merger ratio distribution, and accretion redshift distribution. We find that higher-level subhalos dominate at progressively lower peak mass in the PMF and are more likely to originate from major mergers than lower-level ones. At a given mass ratio at accretion time, the subhalo accretion rates at each level track the growth rate of the host halo. At a fixed final mass ratio, however, the accretion redshift distribution of subhalos depends on the subhalo level, peak mass, and host mass. Higher-level and higher-mass-ratio subhalos tend to be accreted more recently, and more massive halos also accrete their subhalos more recently. Our model provides a concise summary of simulation results and can serve as a basis for further theoretical understanding of the hierarchical structure formation.
@@ 1 Introduction
The existence of dark matter subhalos is a robust prediction of the hierarchical structure formation paradigm. These objects are the self-bound substructures residing in a larger halo as remnants from previous halo mergers. Hence, the properties and distributions of subhalos carry important information about the hierarchical structure formation process. As galaxies form and migrate with halos, subhalos then correspond to the hosts of satellite galaxies inside a larger galaxy system. Understanding the evolution and distribution of subhalos is thus crucial for studying many astrophysical problems, ranging from the structure and origin of stellar streams and the stellar halo (e.g., Banik et al., 2018; Bovy, 2016; Vazquez Segovia et al., 2023; Erkal et al., 2016; He et al., 2024b; Tan et al., 2024), the formation of globular clusters (e.g., Madau et al., 2020; Carlberg & Keating, 2022; Vitral & Boldrini, 2022) to the cosmological formation and evolution of galaxies (e.g., Springel et al., 2001; Yang et al., 2012). The properties of subhalos may also provide sensitive probes to the particle nature of dark matter (e.g., Nadler et al., 2021; Buckley & Peter, 2018; He et al., 2023) and their distributions could also have important consequences on the indirect detection signal (e.g., Springel et al., 2008; Gao et al., 2011; Han et al., 2012a, 2016).
@@
One of the key areas of research in this field is to clarify how dark matter subhalos occupy host halos, how their orbits evolve within these halos, and how they interact and exchange mass with the host halo. The starting point for understanding these questions lies in the initial conditions for subhalo formation, including the initial orbit distribution (Benson, 2005; Wetzel, 2011; Li et al., 2020) and the halo merger rate (e.g. Lacey & Cole, 1993; Fakhouri & Ma, 2008a; Stewart et al., 2009; Genel et al., 2009; Neistein & Dekel, 2008; Poole et al., 2017). Many works have attempted to establish empirical laws for the merger rate in numerical simulations. It is generally found that the merger rate takes on a power law form at the low mass ratio end with an exponential decay at the high ratio end (Fakhouri & Ma, 2008b, 2009; Fakhouri et al., 2010; Stewart et al., 2009; Genel et al., 2009). However, different works have reported different dependence of the merger rate on halo mass and redshift. For example, Fakhouri & Ma (2008b); Fakhouri et al. (2010) found a weak dependence on the progenitor mass and redshift using the Millennium simulation Springel et al. (2001) up to $z=16$. On the other hand, Stewart et al. (2009); Genel et al. (2009) found a more rapid increase with respect to progenitor mass and redshift for $z$ smaller than 1, and Poole et al. (2017) reported a slightly higher major merger rate with a slower increase of merger rate across redshift. The differences among various works may arise from both complexities in the merger tree construction and differences in the background cosmologies. For the former, it is widely accepted that the performances of halo merger trees can be sensitive to the choice of halo finders and the methods to handle pathological cases in merger trees (Srisawat et al., 2013a; Avila et al., 2014; Behroozi et al., 2015a; Knebe et al., 2013). For example, Fakhouri & Ma (2008b) showed that different algorithms for addressing the ’fragmentation’ problem had a pronounced effect on the merger rate. For the latter, different $\Omega_{\mathrm{M_{0}}}$ and $\sigma_{8}$ lead to different growth rates and variances of the smoothed density field separately, thus affecting the halo merger rate globally (Lacey & Cole, 1994; Amoura et al., 2024).
@@
In response to these challenges,  Dong et al. (2022) demonstrated that if the merger rate per halo was normalized by the logarithmic mass growth rate of host halos, the newly defined specific merger rate, expressed as $\mathrm{d}N_{\mathrm{merge}}/N_{\mathrm{halo}}/\mathrm{d}\xi/\mathrm{dln}M_{h}$, exhibits a remarkable universality across the redshift range of $[0,5]$ and the halo mass range of $[\mathrm{10^{12},10^{14}}]\mathrm{M_{\odot}}$. This universality can also be maintained under different background cosmologies with varying values of $\Omega_{\mathrm{M_{0}}}$ and $\sigma_{8}$.  Dong et al. (2022) also pointed out that the specific merger rate is equivalent to the derivative of the PMF  (van den Bosch et al., 2005; Giocoli et al., 2008b) — an alternative approach in describing the distribution of the merging population, which specifies the number of progenitor halos of a given infall mass accumulated through all previous redshifts. Thus, it is expected that the PMF should exhibit the same universality as the merger rate.
In fact, the PMF has been extensively studied in simulations (e.g. Giocoli et al., 2008c; Li & Mo, 2009; Han et al., 2018), Monte-Carlo merger trees (e.g. van den Bosch et al., 2005; Jiang & van den Bosch, 2014), and theoretical models (e.g. Salvador-Solé et al., 2022; Giocoli et al., 2008a; Yang et al., 2011). All of them have demonstrated that the PMF can be largely described by a Schechter function which is independent of host halo mass. Relatively fewer studies, however, have systematically investigated the universality of the PMF across redshift in cosmological simulations (Li & Mo, 2009; Giocoli, 2010), and there have been no simulation investigations of it for varying cosmologies. Therefore, it remains to be tested how well the PMF is universal across mass, redshift, and cosmology, in correspondence to the universality of the specific halo merger rate.
@@
In this study, we systematically examine the universality of the PMF in simulations, addressing several significant issues that arise in this context. First of all, a subhalo finder and a merger tree builder are needed to construct the subhalo population and their evolution histories, the choice of which affects the quality of the derived PMF. Secondly, a definition for the membership of the subhalo, as well as a definition for the peak mass, are required when constructing the PMF. Lastly, depending on the halo finder and mass definition, the masses of the subhalos may depend on the resolution of the simulation. To this end, we adopt hbt+ (Han et al., 2012b, 2018), a time-domain subhalo finder known for its robustness and consistency, to identify subhalos and build the merger tree. We optimize over different combinations of subhalo mass and membership definitions to maximize the level of universality. A number of N-body simulations with different cosmologies and resolutions are employed to investigate the physical and numerical dependence of the results.
@@
A key prediction of the hierarchical structure formation paradigm is that halos merge hierarchically, leading to a tree-like structure describing the merger histories of halos, and a corresponding hierarchy of subhalos with different levels. Previous studies on the merger rate of halos mostly focus on direct mergers among halos, corresponding to the progenitors of first-level subhalos. The use of hbt+allows us to obtain the PMFs for subhalos of different levels. If the PMF of the first level is universal across mass and redshift, an immediate implication is that the PMFs of all levels can be derived through convolutions of the first level function. The resulting universal PMFs are then a direct manifestation of the self-similarity of the halo merger tree, such that the tree truncated at any branch has the same branching properties as the root branch. In the second part of this study, we study how well different levels of the PMF can be modeled this way. The derived model is further used to predict a few population properties of subhalo progenitors in redshift, mass, and level.
@@ 2 Simulation and subhalo catalog
represents the mass ratio between the peak mass of a subhalo and the current mass of the host halo, and $N_{\mathrm{sub}}$ denotes the number of subhalos in the mass bin $[\mu,\mu+\mathrm{d}\mu]$ per host halo. The level of universality in the PMF could differ depending on the exact definitions of the subhalo and host halo masses, as well as on how the subhalos are selected when counting their numbers for each halo. In hbt+, the mass of a subhalo is defined to be its self-bound mass, and we define the peak mass, $m_{\rm peak}$, of a subhalo as the maximum bound mass reached during its evolutionary history. A significant fraction of subhalos could become unresolved (or “disrupted") in simulations due to their high mass loss rates and the finite resolution of the simulation (Han et al., 2016; He et al., 2024a). Additionally, some subhalos could have merged with other subhalos or the host halo (defined as “sink" in hbt+). It is necessary to include the progenitors of these disrupted or merged populations when computing the PMF, in order to account for all the subhalos that have ever merged into the host. In hbt+, a subhalo is tracked continuously with its most-bound particle once it becomes unresolved. This allows us to easily collect all the subhalos that have been accreted into a given host halo, no matter whether these subhalos have become unresolved or not. For the main part of this work, we count all the subhalos, including disrupted or merged ones, in each FoF group when computing the PMF, and define the mass of the host halo as the total bound mass of their member subhalos accordingly. This choice is different from many previous works, which use the virial mass for the host halo mass and only count subhalos inside the virial radius. We discuss the influence of different mass and membership definitions in Section 5.
@@ 3.2 Fitting the PMFs
Figure 1: The level-1 PMF across host halo mass and redshifts in the L600 simulation, as labeled. The gray solid line shows the best-fit double Schechter function at $z=0$, reproduced in all the panels. The light gray shaded region shows a ±$\pm 10\%$ deviation, while the darker shaded region indicates a ±$\pm 5\%$ deviation from the fit. The bottom sub figure in each panel shows the ratio between the data and the fit.
@@
This best-fitting model is overplotted in all the other panels, with a gray-shaded region representing the ±$\pm 10\%$ deviation. No significant variations (>10%) are found between the data points and the fit, except for some data points with large Poisson errors.
Figure 2: The PMFs of different levels and redshifts in the L600 simulation. Subhalos residing in different halo mass bins between $[10^{11},10^{15}]\mathrm{M_{\odot}}$ have been combined in the measurement. For clarity, we have not shown the measurements of the level-1 PMF. The dashed lines of different colors show the best-fit models at $z=0$ according to Equation (4), reproduced in all three panels. Note the parameters controlling $g_{1}$ are fixed as in Fig. 1, while only the $\beta$ parameter is varied to fit the data at other levels. The shaded regions show the ±$\pm 10\%$.
With a model of $g_{1}$ established, we further fit the PMFs of higher levels by varying the only remaining parameter $\beta$. At $z=0$, we obtain a best-fit parameter $\beta=0.726$. As shown in Fig. 2, the model can well match the measurements at different levels simultaneously. The same model is also compared against measurements at other redshifts up to $z=2$, showing generally good agreement. At the highest level, there is a tendency that the PMF becomes underestimated at higher redshifts. This is indicative of a weak redshift evolution in the subhalo PMF. The deviation from the universal model is too small to be noticed in low-level PMFs but becomes amplified in high-level PMFs according to Equation (5). Nevertheless, the PMF at high redshift also becomes poorly resolved in the simulation. The weak redshift evolution will be examined in more detail in an upcoming work using the Jiutian simulation of a higher resolution (Han et al., in prep.). As the level-4 PMF contributes little to the total PMF, the universal model still works very well in predicting the total PMF.

### v3

@@ Abstract
In the $\Lambda$CDM universe, structure formation is generally not a self-similar process, while some self-similarity remains in certain statistics which can greatly simplify our description and understanding of the cosmic structures. In this work, we show that the merger tree of dark matter halos is approximately self-similar by investigating the universality of the subhalo peak mass function (PMF) describing the mass distribution of progenitor halos. Using a set of cosmological simulations and identifying subhalos of different merger levels with hbt+, we verify that the level-1 subhalo PMF is close to universal across halo mass, redshift, and cosmology. This approximate self-similarity allows us to analytically derive the subhalo PMF for subhalos accreted at any level (i.e., for sub-sub…halos) through self-convolutions of the level-1 PMF, and the resulting model shows good agreement with simulation measurements. We further derive a number of analytical properties on the hierarchical origin of subhalos. We show that higher-level subhalos dominate at progressively lower peak mass in the PMF and are more likely to originate from major mergers than lower-level ones. At a given merger mass ratio, the subhalo accretion rates at each level track the growth rate of the host halo. At a fixed final mass ratio, however, subhalos of higher-level, higher-mass-ratio and in more massive haloes tend to be accreted more recently. Matching subhalo peak mass to galaxy mass, these results have direct implications on the hierarchical origin of satellite galaxies.
@@ 1 Introduction
The seek for self-similarity used to be a fundamental topic in the study of structure formation (e.g., Press & Schechter, 1974; Peebles, 1974, 1985; Miyoshi & Kihara, 1975; Davis & Peebles, 1977; Gott & Turner, 1977; Efstathiou et al., 1979; Efstathiou & Eastwood, 1981; Fry & Peebles, 1980; Efstathiou et al., 1988). As gravity is scale-free, when further coupled with a self-similar initial condition, then the process and outcome of structure formation can be expected to be self-similar, so that the distribution and evolution of structures of different sizes follow the same laws when rescaled by their sizes. For example, the collapse of a dark matter halo is self-similar for power-law initial conditions in the Einstein de Sitter (EdS) universe (Fillmore & Goldreich, 1984). The formation of large scale structure is self-similar in the EdS universe with power-law initial conditions (e.g., Efstathiou et al., 1988).
In the $\Lambda$CDM universe, however, structure formation is not exactly self-similar, as the initial power spectrum is scale dependent and dark energy introduces another scale. Despite this, some (approximate) self-similarity are still found in many statistics of the $\Lambda$CDM universe. The identification of these laws can greatly simplify our description of structure formation, while also bringing crucial insights about the physical processes. On one hand, accurate universal models of the clustering of matter in the nonlinear regime have been developed (Hamilton et al., 1991; Peacock & Dodds, 1994; Smith et al., 2003), which started from analytical insights brought by self-similarity (Peebles, 1974, 1980; Davis & Peebles, 1977). On the other hand, the evolution of the cosmic structure described in terms of dark matter halos has also revealed much self-similarity. For example, the Extended Press Schechter theory predicts that the halo mass function, as well as its time evolution, is universal when formulated with abstract variables such as the peak height, $\nu$, and collapse barrier, $\omega$, which can be regarded as a manifestation of self-similarity in this particular parameter space (e.g., Press & Schechter, 1974; Lacey & Cole, 1993, 1994; Sheth et al., 2001; Moreno et al., 2008; Neistein et al., 2010). However, in terms of intuitive properties such as the mass of halos, self-similarity becomes rarer to find, due to the complexities introduced by scale dependence in the initial condition and structure formation. A notable exception is found in the distribution of subhalos. The subhalo mass function, expressed as a function of the ratio between the subhalo mass and the host halo mass, has long been shown to be largely self-similar (e.g., Gao et al., 2004; Stewart et al., 2008; Giocoli et al., 2008b; Jiang & van den Bosch, 2014). Recently, Yang & Yu (2023) also observed self-similarity in the graph showing the spatial hierarchy of the surviving subhalos in the Fire and Fire2 simulation.
The origin of the approximate self-similarity in the subhalo distribution has to be traced back to how halos merge and their tidal evolutions thereafter (e.g., Han et al., 2016; He et al., 2024a). The distribution of the merger events can be studied in two equivalent approaches, by either studying the merger rate of halos, or by studying the distribution of progenitor masses. It thus can be generally expected that some self-similarity could be found in these two statistics.
Along the first line, it has taken some efforts before self-similarity is identified in numerical simulations. It has been generally found that the merger rate takes on a power law form at the low mass ratio end with an exponential decay at the high ratio end (Fakhouri & Ma, 2008a, 2009; Fakhouri et al., 2010; Stewart et al., 2009; Genel et al., 2009). However, different works have reported different dependencies of the merger rate on halo mass and redshift. For example, Fakhouri & Ma (2008a) found a weak dependence on the progenitor mass ($\propto M_{\rm h}^{0.08}$) and redshift ($\propto\frac{d\delta_{c}}{dz}^{0.3}$) using the Millennium simulation Springel et al. (2001) up to $z=16$. On the other hand, Stewart et al. (2009); Genel et al. (2009) found a more rapid increase with respect to progenitor mass and redshift for $z$ smaller than 1 ($\propto M_{\rm h}^{0.15}(\frac{d\delta_{c}}{dz})^{2}$;$\propto M_{\rm h}^{0.12}(\frac{d\delta_{c}}{dz})$ separately), and Poole et al. (2017) reported a slightly higher major merger rate with a slower increase of the merger rate across redshift.
@@
The differences among various works may arise from both complexities in the merger tree construction and differences in the background cosmologies. For the former, it is widely accepted that the performances of halo merger trees can be sensitive to the choice of halo finders and the methods to handle pathological cases in merger trees (Srisawat et al., 2013a; Avila et al., 2014; Behroozi et al., 2015a; Knebe et al., 2013). For example, Fakhouri & Ma (2008a) showed that different algorithms for addressing the ’fragmentation’ problem had a pronounced effect on the merger rate. For the latter, different $\Omega_{\mathrm{M_{0}}}$ and $\sigma_{8}$ lead to different growth rates and variances of the smoothed density field separately, thus affecting the halo merger rate globally (Lacey & Cole, 1994; Amoura et al., 2024).
@@
In a more recent work,  Dong et al. (2022) demonstrated that if the merger rate per halo was normalized by the logarithmic mass growth rate of host halos, the newly defined specific merger rate, expressed as $\mathrm{d}N_{\mathrm{merge}}/N_{\mathrm{halo}}/\mathrm{d}\xi/\mathrm{dln}M_{h}$, exhibits a remarkable universality across the redshift range of $[0,5]$ and the halo mass range of $[\mathrm{10^{12},10^{14}}]\mathrm{M_{\odot}}$. This universality can also be maintained under different background cosmologies with varying values of $\Omega_{\mathrm{M_{0}}}$ and $\sigma_{8}$.  Dong et al. (2022) also pointed out that the specific merger rate is equivalent to the derivative of the Peak Mass Function (PMF) — an alternative approach in describing the distribution of the merging population, which specifies the number of progenitor halos of a given infall mass accumulated through all previous redshifts. Thus, it is expected that the PMF should exhibit the same universality as the merger rate.
The PMF has been extensively studied using various approaches, including numerical simulations (e.g., Giocoli et al., 2008b; Stewart et al., 2008; Li & Mo, 2009; Han et al., 2018; Xie & Gao, 2015; Chua et al., 2017), Monte Carlo merger trees (e.g., van den Bosch et al., 2005; Jiang & van den Bosch, 2014), and theoretical models (e.g., Salvador-Solé et al., 2022; Giocoli et al., 2008a; Yang et al., 2011). These studies have consistently shown that the PMF can be well described by a Schechter function. While some studies suggest the PMF is universal across halo mass ranges (e.g., Giocoli et al., 2008b; Li & Mo, 2009; Han et al., 2018), others, such as Xie & Gao (2015), report a clear mass dependence in the Aquarius and Phoenix simulations, particularly when excluding ‘re-accreted’ and ‘ejected’ subhalos from the unevolved population. In contrast, the redshift dependence of the PMF has been less explored, with only a few studies examining its universality in cosmological simulations (e.g., Li & Mo, 2009; Giocoli, 2010). Furthermore, the influence of varying cosmological parameters on the PMF remains largely unstudied in simulations. Therefore, it remains to be investigated whether and to what extent the PMF can be modeled as universal, and what subhalo mass and membership definitions lead to the maximum amount of universality across halo mass, redshift, and cosmology, in correspondence to the universality of the specific halo merger rate.
@@
In this study, we adopt HBT+ subhalo finder (Han et al., 2012b, 2018), a time-domain subhalo finder known for its robustness and consistency, to identify subhalos and build the merger tree. We systematically examine the universality of the PMF in simulations, addressing several significant issues that arise in this context. Firstly, a definition for the membership of the subhalo, as well as a definition for the peak mass, are required when constructing the PMF. Second, depending on the halo finder and mass definition, the masses of the subhalos may depend on the resolution of the simulation. To this end, we investigate different combinations of subhalo mass and membership definitions, and identify the one that best preserves the universality of the PMF. A number of N-body simulations with different cosmologies and resolutions are employed to investigate the physical and numerical dependence of the results.
@@
A key prediction of the hierarchical structure formation paradigm is that halos merge hierarchically, leading to a tree-like structure describing the merger histories of halos, and a corresponding hierarchy of subhalos with different levels. Previous studies on the merger rate of halos mostly focus on direct mergers among halos, corresponding to the progenitors of first-level subhalos (e.g., Fakhouri & Ma, 2008b, 2009; Fakhouri et al., 2010; Genel et al., 2009; Poole et al., 2017; Dong et al., 2022). In terms of the PMF, subhalo progenitors without selection in level, i.e., at all levels as a whole, are also frequently studied (e.g., Li & Mo, 2009; Xie & Gao, 2015; Chua et al., 2017; Han et al., 2018). The use of hbt+allows us to obtain the PMFs for subhalos at different levels separately. If the PMF of the first level is universal across mass and redshift, an immediate implication is that the PMFs of all levels can be derived through convolutions of the first level function. The resulting universal PMFs are then a direct manifestation of the self-similarity of the halo merger tree, such that the tree truncated at any branch has the same branching properties as the root branch. In the second part of this study, we study how well different levels of the PMF can be modeled this way. The derived model is further used to predict a few population properties of subhalo progenitors in redshift, mass, and level.
As subhalos correspond to the hosts of satellite galaxies inside a larger galaxy system, the universal and analytical model for the multi-level PMFs obtained here may find applications in a wide range of studies related to subhalos, including the structure and origin of stellar streams and the stellar halo (e.g., Banik et al., 2018; Bovy, 2016; Vazquez Segovia et al., 2023; Erkal et al., 2016; He et al., 2024b; Tan et al., 2024), the formation of globular clusters (e.g., Madau et al., 2020; Carlberg & Keating, 2022; Vitral & Boldrini, 2022), the cosmological formation and evolution of galaxies (e.g., Springel et al., 2001; Yang et al., 2012), and the particle nature of dark matter (e.g., Nadler et al., 2021; Buckley & Peter, 2018; He et al., 2023; Springel et al., 2008; Gao et al., 2011; Han et al., 2012a, 2016).
@@ 2 Simulation and subhalo catalog
represents the mass ratio between the peak mass of a subhalo and the current mass of the host halo, and $N_{\mathrm{sub}}$ denotes the number of subhalos in the mass bin $[\mu,\mu+\mathrm{d}\mu]$ per host halo. The level of universality in the PMF could differ depending on the exact definitions of the subhalo and host halo masses, as well as on how the subhalos are selected when counting their numbers for each halo. In hbt+, the mass of a subhalo is defined to be its self-bound mass, and we define the peak mass, $m_{\rm peak}$, of a subhalo as the maximum bound mass reached during its evolutionary history. A significant fraction of subhalos could become unresolved (or “disrupted") in simulations due to their high mass loss rates and the finite resolution of the simulation (Han et al., 2016; He et al., 2024a). Additionally, some subhalos could have merged with other subhalos or the host halo (defined as “sink" in hbt+). It is necessary to include the progenitors of these disrupted or merged populations when computing the PMF, in order to account for all the subhalos that have ever merged into the host. In hbt+, a subhalo is tracked continuously with its most-bound particle once it becomes unresolved. This allows us to easily collect all the subhalos that have been accreted into a given host halo, no matter whether these subhalos have become unresolved or not. For the main part of this work, we count all the subhalos, including disrupted or merged ones, in each FoF group when computing the PMF, and define the mass of the host halo as the sum of the bound masses of their member subhalos accordingly. This choice is different from many previous works, which use the virial mass for the host halo mass and only count subhalos inside the virial radius. We discuss the influence of different mass and membership definitions in Section 5. As the PMF at each redshift is determined independently from the subhalos belonging to the host halo at that time, any changes of subhalo membership including fly-by or ejection of subhalos are automatically accounted for.
@@ 3.2 Fitting the PMFs
Figure 1: The level-1 PMF across host halo mass and redshifts in the L600 simulation, as labeled. The data points show the average PMF over all halos in each specific mass bin and redshift. The gray solid line shows the best-fit double Schechter function at $z=0$, reproduced in all the panels. The light gray shaded region shows a ±$\pm 10\%$ deviation, while the darker shaded region indicates a ±$\pm 5\%$ deviation from the fit. The bottom sub figure in each panel shows the ratio between the data and the fit. The errorbars show the uncertainty on the average PMF estimated from the Poisson noise in the total subhalo count in each bin.
@@
This best-fitting model is overplotted in all the other panels, with a gray-shaded region representing the ±$\pm 10\%$ deviation. No significant variations (>10%) are found between the data points and the fit, except for some data points with large Poisson errors. We also conduct a brief analysis on the halo-to-halo variation on $g_{1}$ in Appendix A. The results show that the scatter of the $g_{1}$ among different halos is close to or smaller than the Poisson noise in the region $\mu>10^{-2}$. For less massive subhalos $\mu<10^{-2}$, the $g_{1}$ shows a mildly broader distribution than the Poisson distribution($1.61.0<\sigma/\sigma_{\rm Poisson}<$ 1.6), indicating the presence of a small amount of halo-to-halo scatter on top of the Poisson noise.
Figure 2: The PMFs of different levels and redshifts in the L600 simulation. Subhalos residing in different halo mass bins between $[10^{11},10^{15}]\mathrm{M_{\odot}}$ have been combined in the measurement. For clarity, we have not shown the measurements of the level-1 PMF. The dashed lines of different colors show the best-fit models at $z=0$ according to Equation (4), reproduced in all three panels. Note the parameters controlling $g_{1}$ are fixed as in Fig. 1, while only the $\beta$ parameter is varied to fit the data at other levels. The shaded regions show the ±$\pm 10\%$. The errorbars show the uncertainty on the average PMF estimated from the Poisson noise in the total subhalo count in each bin.
With a model of $g_{1}$ established, we further fit the PMFs of higher levels by varying the only remaining parameter $\beta$. At $z=0$, we obtain a best-fit parameter $\beta=0.726$. As shown in Fig. 2, the model can well match the measurements at different levels simultaneously. The same model is also compared against measurements at other redshifts up to $z=2$, showing generally good agreement. At the highest level, there is a tendency that the PMF becomes underestimated at higher redshifts. This is indicative of a weak redshift evolution in the subhalo PMF. The deviation from the universal model is too small to be noticed in low-level PMFs but becomes amplified in high-level PMFs according to Equation (5). Nevertheless, the PMF at high redshift also becomes poorly resolved in the simulation. The weak redshift evolution will be examined in more detail in an upcoming work using the Jiutian simulation of a higher resolution Han et al. (2025). As the level-4 PMF contributes little to the total PMF, the universal model still works very well in predicting the total PMF.